Boolean logic is the form of algebra in which all values are needed to either TRUE or FALSE normally through AND,OR and NOT operator.
a)How do the number of result change upon using an "AND" operator and an "OR" operator?EXPLAIN your answer with the help of relevant examples.
b)What is the name for the combination logic gate of an "AND" operator followed by a "NOT" operator?Explain the whole operation and provide the reasons on why do we need this combination.

Answer by Edwin McCravy(20055)

If AND is between two sentences, the result will be FALSE
unless BOTH the sentence on the left of AND and the sentence
on the right of AND are TRUE. That's the only situation
that the result will be TRUE. 

AND is more often FALSE than it is TRUE.

AND requires both sides to be TRUE for the result to be TRUE,
otherwise it's FALSE.

------------------------------------------

If OR is between two sentences, the result will be TRUE
unless BOTH the sentence on the left of AND and the sentence
on the right of AND are FALSE. That's the only situation
that the result will be FALSE.

OR is more often TRUE than it is FALSE.

OR requires both sides to be FALSE for the result to be FALSE, 
otherwise it's TRUE.

------------------------------------------

Examples:

8+1=11 AND 6x2=15

That is FALSE because both sides of AND are FALSE

-------------------------------

2+3=5 AND 5-1=3    

That is FALSE because both sides of AND are not TRUE,
even though the left side is TRUE, the right side is not,
so the result is FALSE.

-------------------------------

2+5=8 AND 6x2=12

That is FALSE because both sides of AND are not TRUE,
even though the right side is TRUE, the left side is not,
so the result is FALSE.

-------------------------------


3+6=9 AND 7-3=4

That is TRUE because both sides of AND are TRUE. Both
sides of AND being TRUE is the only time the result
will be TRUE.

-------------------------------
-------------------------------

-------------------------------

2+3=5 OR 5-1=3    

That is TRUE because both sides of OR are not FALSE,
even though the right side is FALSE, the left side is TRUE,
so the result is TRUE.

-------------------------------

2+5=8 OR 6x2=12

That is TRUE because both sides of AND are not FALSE,
even though the left side is FALSE, the right side is TRUE.
so the result is TRUE.

-------------------------------

3+6=9 OR 7-3=4

That is TRUE because both sides of OR are TRUE.

-------------------------------
-------------------------------

An easy way to think of it is

OR needs only ONE side of it to be TRUE in order for the result to be TRUE.

but

AND needs BOTH sides of it to be TRUE in order for the result to be TRUE.

Edwin